# SAT-based Exact Modulo Scheduling Mapping for Resource-Constrained CGRAs

### v1

@@ 1. Introduction
While Application Specific Integrated Circuits (ASICs) accelerators have been largely adopted in these scenarios due to their efficiency, they are limited by their fixed functionality and their lengthy and costly design time. The reconfigurability of Field Programmable Gate Arrays (FPGAs) can address these limitations, but at the cost of lower power efficiency compared to ASICs, due to their fine-grained structure.
@@
Coarse-Grain Reconfigurable Arrays (CGRAs) provide a meet-in-the-middle approach, providing run-time reconfigurability at the level of arithmetic operations, with very high computational efficiency (Li et al., 2021; Karunaratne et al., 2017). These characteristics are well suited in domains such as streaming and multimedia applications (Li et al., 2021; Akbari et al., 2018; Karunaratne et al., 2017; Oh et al., 2009; Lee et al., 2015; Wijerathne et al., 2019; Duch et al., 2017), but also in the edge domain, where they have been shown to enable flexible hardware acceleration in resource-constrained scenarios.
A CGRA is a mesh of Processing Elements (PEs) organized in a two-dimensional grid; each PE contains an Arithmetic-Logic Unit (ALU) and a number of internal registers. PEs are connected to their neighbors according to a mesh topology. In addition, shared connections are usually present from PEs to external memory to load inputs and store outputs. Computation in CGRAs is organized in CGRA-cycles, during which every PE performs an operation. The list of supported operations defines the Instruction Set Architecture (ISA) of a CGRA.
@@
One of the fundamental challenges of CGRA exploitation is the compilation process, i.e. the translation of high-level source code onto CGRA code, while taking advantage of the parallelism offered by the architecture. To do so, a technique called modulo scheduling is traditionally employed, which constructs a Data Flow Graph (DFG) from a Compute-Intensive Loop (CIL), and then maps DFG nodes onto architecture PEs in an interleaved manner so that nodes from different iterations coexists in the same CGRA-cycle, minimizing the overall latency of each iteration. Figure 1 portrays an overview of this process, where a section of C code is extracted and mapped into a $2\times 2$ CGRA. First, a loop is marked for CGRA acceleration. Then, its DFG, which illustrates operations as nodes and data dependencies as edges, is derived. The DFG is then scheduled over the PEs of the target architecture, at different CGRA-cycles, while abiding to resource and dependency constraints. A possible modulo scheduling of the simple DFG in Figure 1.b is shown in Figure 1.e, highlighting the interleaving among loop iterations.
@@
As the number of PEs (and the resources available to them) is reduced to fit on edge devices, mapping tools are faced with growing constraints. Existing techniques for modulo scheduling, described in Section 2, rely mainly on heuristics to schedule, place, and then route operations and data on the PEs. Here, we propose to address the mapping problem using SAT-MapIt, a satisfiability (SAT)-based formulation where data dependency, architectural constraints, and schedule are expressed as Boolean constraints. A valid mapping is then identified by determining if an assignment of the variables exists to satisfy the constructed Boolean formula. We show that this technique is able to efficiently explore the space of possible mappings better than State of the Art (SoA) techniques, hence producing high-performance mappings even for a tightly constrained CGRA topology.
@@
We showcase the performance of our approach as part of a complete code-to-hardware framework. The flow starts from annotated C code and has at its endpoint a configuration bitstream that governs the execution of a (simulated or synthesized) CGRA. To this end, we herein target the open-hardware CGRA design in (Rodríguez Álvarez et al., 2023). We highlight that compiler-level metrics (such as Iteration Interval (II) and Utilization (U)) do indeed correlate with the run-time performance of a mapped CIL. However, hardware analysis (and hence the availability of a link between the compiler and the hardware) is the key to accurately assess the energy and timing requirements for a given CIL and CGRA architecture.
@@ 2. Related Work
A second category of approaches addresses the mapping problem with Integer Linear Programming (ILP) or Boolean Satisfiability formulations. In (Chin and Anderson, 2018) the authors propose an ILP formulation approach and prove the feasibility of mapping in the given number of CGRA-cycles. Similarly, (Miyasaka et al., 2020) propose to use a SAT solver instead of an ILP solver to identify a valid solution. These works represent a first effort towards exact formulations, such as ours, but they are not capable of achieving a modulo scheduled solution. Our SAT formulation, initially introduced in (Tirelli et al., 2023), is to the best of our knowledge the first to propose an exact solution to the modulo scheduling problem, and yet scale to DFG sizes that were previously only tackled via heuristics.
@@ 3.1. Compilation
To accelerate an application to a CGRA, a CIL is identified in the application, as in Figure 2.a; the identification can either be performed automatically through techniques such as, for example, (Zacharopoulos et al., 2018), or manually by the programmer through pragma annotations, as done in this work. Then, CIL must be compiled, in order to be translated into CGRA instructions. The first step in this process, depicted in Figure 2, is to generate a semantically equivalent version in Intermediate Representation (IR) (LLVM IR is our chosen one) and from there to DFG. DFGs are directed graphs in which nodes represent operations, edges represent dependency relations between operations, and loop-carried dependencies correspond to back-edges.
@@ 3.2. Modulo Scheduling
Modulo scheduling is a compilation technique that enables efficient execution of a CIL, by executing multiple iterations of it in an interleaved manner. As exemplified in 2(a), a modulo-scheduled CIL is divided into three stages: prologue, kernel, and epilogue. Prologue and epilogue are one-time executed stages: the former is used to prepare the data to feed the pipeline, the latter to reorganize them at the end. The kernel is instead the steady state that executes multiple times and includes the operations to be parallelized through pipelining. The goal of modulo scheduling is to pipeline as effectively as possible the execution of CIL operations, and this corresponds to minimizing the II, i.e. the duration of the kernel stage – 3 CGRA-cycles in our example. Therefore, the mapping problem consists in finding a legal modulo schedule for a CIL, performing the placement and routing of operations in a constrained 3D space represented by the PEs dimensions and by time. An example of legal mapping for the DFG in our running example is shown in Figure 3.b.
@@ 4.1. Schedule creation
Table 2. CIL Mobility Schedule creation.
In blue iteration 0, in green iteration 1
MS
Time Nodes
0 1 2 3 4
1 1 2 4 5 7 10
2 1 2 6 7 10 11
3 2 8 10 11
4 9 11
@@ 4.2. SAT formulation
Equation 4 will be used on each node of the DFG and each $\beta$ generated will be added to the SAT solver. In our example with $n_{3}$ Equation 4 becomes:
@@ 4.3. Register Allocation
In our CFN formulation, to have a more straightforward set of constraints, we chose not to include information about which register should store the data produced by the node on every PE. We hence treat the RA task as a separate problem.
After the solver returns a mapping, the RA phase needs to ensure that the CGRA has an adequate number of registers to execute the CIL correctly. SAT-MapIt  as a separate phase subsequent to SAT solving, exploits the SSA format of the input code and looks for an optimal solution(Hack et al., 2006) of the RA problem.
For each PE in the CGRA, we construct an interference graph that we then attempt to color using $n$ colors, where $n$ is the number of registers available on each PE. If the coloring succeeds, we know that no further action is needed, and the toolchain provides the assembly code for the CGRA. However, if the coloring process fails, we increment the II and initiate a new search. While it’s possible to explore additional graph splitting techniques to make the graph colorable, we do not consider them herein, as such optimizations are orthogonal to the mapping problem we address in this work.
@@ 5. Experimental Results: SAT-MapIt solutions
In this section we evaluate the effectiveness of SAT-MapIt on a set of CILs from MiBench and Rodinia benchmark suites, shown in Table 3. We compare the obtained II, and the time to find it, with respect to two techniques of the SoA: RAMP(Dave et al., 2018) and PathSeeker (Balasubramanian and Shrivastava, 2022).
@@ 5.1. Experimental Setup
Table 3. List of benchmarks with graph information
Suite CIL #nodes #edges
MiBench sha 30 33
sha2 26 28
gsm 20 24
patricia 42 46
bitcount 26 29
basicmath 19 20
stringsearch 16 16
Rodinia backpropagation 35 39
nw 16 16
srand 22 22
hotspot 67 76
@@ 5.3. SAT-MapIt uses tight resources better
5.4. SAT-MapIt is faster when run-times are high
Given that the II found are better than SoA, it may be interesting to also analyze the time needed to find such solutions; this is reported in Table 4. The following statements can be made when looking at these numbers. 1) Our tools running time is longer than SoA in 26 out of 44 experiments. 2) However, in these 26 cases the average time difference is only 15.28 seconds, with a standard deviation of $4.973$) On the other hand, in the 18 cases in which our tool is faster, the average time difference is 962.24 seconds, with a standard deviation of $1438.78$. This shows that SAT-MapIt is significantly faster when it matters, i.e. when computation times are high.
@@ 5.5. Limitations of SAT-MapIt
Currently, our tool does not apply any routing strategies. This limitation manifests in the sha CIL of a $5\times 5$ CGRA, where we achieve an II of 3, while (Dave et al., 2018) and (Balasubramanian and Shrivastava, 2022) can find an II of 2 by adding a routing node. This is the only case, out of the 44 experiments shown, where the effect of this limitation can be noticed. Implementing routing capabilities in SAT-MapIt will be the subject of future work.
@@ 7.1. Experimental Setup
We consider square openEdgeCGRA instances with sizes $2\times 2$, $3\times 3$ and $4\times 4$, referred to as $D2$, $D3$ and $D4$, respectively, in the following. Benchmark CILs from the MiBench suite were executed on the architectures. To compare with the SoA in Section 5.2 we used the same DFGs generated by the RAMP toolchain; however those DFGs are not inherently compatible with the target CGRA’s ISA, preventing the direct execution of experiments using identical graphs. To address this discrepancy, we recalculated the DFGs for each CIL to align with the instruction set of openEdgeCGRA, for this part of the experiments. Consequently, this resulted in different II values compared to Figure 7.

### v2

@@ 1. Introduction
While Application Specific Integrated Circuits (ASICs) accelerators have been largely adopted in these scenarios due to their efficiency, they are limited by their fixed functionality and their lengthy and costly design time. The reconfigurability of Field Programmable Gate Arrays (FPGAs) can address these limitations, but at the cost of lower power and area efficiency, because, due to their fine-grained structure, the hardware overhead required to reconfigure an FPGA can add power and area overheads of up to 10×10$\times$10 × and 20×20$\times$20 ×, respectively (Lin et al., 2014; Kuon and Rose, 2006).
@@
Coarse-Grain Reconfigurable Arrays (CGRAs) provide a meet-in-the-middle approach, providing run-time reconfigurability at the level of arithmetic operations, with very high computational efficiency (Li et al., 2021; Karunaratne et al., 2017). These characteristics are well suited in domains such as streaming and multimedia applications (Li et al., 2021; Akbari et al., 2018; Karunaratne et al., 2017; Oh et al., 2009; Lee et al., 2015; Wijerathne et al., 2019; Duch et al., 2017a), but also in the edge domain, where they have been shown to enable flexible hardware acceleration in resource-constrained scenarios.
[…]
While large CGRAs meshes can be the preferred choice in high-performance architectures, edge scenarios have to cope with tight power and area budgets. In this cases, small, time-multiplexed CGRAs constitute a promising avenue towards the design of versatile and energy-efficient accelerators. Indeed, important energy and latency savings with respect to low-power CPU execution have been reported in the State of the Art (SoA) from employing resource-constrained CGRAs. In  (Duch et al., 2017b) a 37% energy decrease is reported when running compressed-sensing algorithms. (Carpentieri et al., 2024) showcases 3.4×3.4$\times$3.4 × and $\times9$ energy and latency improvements, respectively, when performing convolutions. The architecture of (Kim et al., 2014) reports energy improvements of almost 60% when compared to a CPU-only platform from the SoA.
@@
One of the fundamental challenges of CGRA exploitation is the compilation process, i.e. the translation of high-level source code onto CGRA code, while taking advantage of the parallelism offered by the architecture. To do so, a technique called modulo scheduling is traditionally employed, which constructs a Data Flow Graph (DFG) from a Compute-Intensive Loop (CIL), and then maps DFG nodes onto architecture PEs in an interleaved manner so that nodes from different iterations coexists in the same CGRA-cycle, minimizing the overall latency of each iteration.
[…]
Figure 1 portrays an overview of this process, where a section of C code is extracted and mapped into a $2\times 2$ CGRA. First, a loop is marked for CGRA acceleration. Then, its DFG, which illustrates operations as nodes and data dependencies as edges, is derived. The DFG is then scheduled over the PEs of the target architecture, at different CGRA-cycles, while abiding to resource and dependency constraints. A possible modulo scheduling of the simple DFG in Figure 1.b is shown in Figure 1.e, highlighting the interleaving among loop iterations.
@@
As the number of PEs (and the resources available to them) is reduced to fit on edge devices, mapping tools are faced with growing constraints. Existing techniques for modulo scheduling, described in Section 2, rely mainly on heuristics to schedule, place, and then route operations and data on the PEs. Here, we propose to address the mapping problem using SAT-MapIt, a satisfiability (SAT)-based formulation where data dependency, architectural constraints, and schedule are expressed as Boolean constraints. A valid mapping is then identified by determining if an assignment of the variables exists to satisfy the constructed Boolean formula. We show that this technique is able to efficiently explore the space of possible mappings better than SoA techniques, hence producing high-performance mappings even for a tightly constrained CGRA topology.
@@
Building upon (Tirelli et al., 2023) we further clarify the toolchain’s architecture, enriching the SAT formulation and the description of the register allocation phase to ensure a deeper understanding of its functionality and efficiency. Additionally, we showcase the performance of our approach as part of a complete code-to-hardware framework, allowing the evaluation of speedups with respect to non-accelerated executions and contrasting mapping figures of merits with run-time metrics. The flow starts from annotated C code and has at its endpoint a configuration instructions list (bitstream) that governs the execution of a (simulated or synthesized) CGRA. To this end, we herein target the open-hardware CGRA design in (Rodríguez Álvarez et al., 2023). We highlight that compiler-level metrics (such as Iteration Interval (II) and Utilization (U)) do indeed correlate with the run-time performance of a mapped CIL. However, hardware analysis (and hence the availability of a link between the compiler and the hardware) is the key to accurately assess the energy and timing requirements for a given CIL and CGRA architecture.
@@ 2. Related Work
A second category of approaches addresses the mapping problem with Integer Linear Programming (ILP) or Boolean Satisfiability formulations. In (Chin and Anderson, 2018) the authors propose an ILP formulation approach and prove the feasibility of mapping in the given number of CGRA-cycles. Similarly, (Miyasaka et al., 2020) propose to use a SAT solver instead of an ILP solver to identify a valid solution.
The SAT formulation proposed in this paper addresses a critical limitation of the approach in [22], by managing back-edges in dataflow graphs and hence being capable of modelling loop-carried dependencies, where the formulation in [22] could only consider a feed-forward, pipelined solution. This is an important advancement both theoretical – because in the absence of back-edges the problem admits a more trivial solution – and practical, as the capability to process loop-carried dependencies results in the ability to accelerate more diverse loops.
Our SAT formulation, initially introduced in (Tirelli et al., 2023), is to the best of our knowledge the first SAT-based solution to the modulo scheduling problem.
@@ 3.1. Compilation
To accelerate an application to a CGRA, a CIL is identified in the application, as in Figure 2.a; the identification can either be performed automatically through techniques such as, for example, (Zacharopoulos et al., 2018), or manually by the programmer through pragma annotations, as done in this work. Then, CIL must be compiled, in order to be translated into CGRA instructions.
Compilation is in general the process of translating a program written in high-level code into a semantically-equivalent one in low-level code. The first step in this process, depicted in Figure 2.b is to generate a version in Intermediate Representation (IR) (LLVM IR is our chosen one) from the original high-level-code.
The second step is to go from there to DFG: DFGs are directed graphs in which nodes represent operations, edges represent dependency relations between operations, and loop-carried dependencies correspond to back-edges. Note that the DFG reflects the number and types of instructions in the LLVM IR, which may differ from those seen in the original high-level code, as an effect of semantic-preserving compiler transformations.
@@ 3.2. Modulo Scheduling
Modulo scheduling is a compilation technique that enables efficient execution of a CIL, by executing multiple iterations of it in an interleaved manner. As exemplified in 3(a), a modulo-scheduled CIL is divided into three stages: prologue, kernel, and epilogue. Prologue and epilogue are one-time executed stages: the former is used to prepare the data to feed the pipeline, the latter to reorganize them at the end. The kernel is instead the steady state that executes multiple times and includes the operations to be parallelized through pipelining. The goal of modulo scheduling is to pipeline as effectively as possible the execution of CIL operations, and this corresponds to minimizing the II, i.e. the duration of the kernel stage – 3 CGRA-cycles in our example. Therefore, the mapping problem consists in finding a legal modulo schedule for a CIL, performing the placement and routing of operations in a constrained 3D space represented by the PEs dimensions and by time. An example of legal mapping for the DFG in our running example is shown in Figure 3.b.
@@ 3.4. SAT problem
Encoding the mapping problem within the framework of a SAT problem offers several distinct benefits. The primary advantage of using SAT solvers in our methodology is to exploit their capability to conduct a more structured and comprehensive exploration of the solution space and, at the same time, find exact solutions. Ensuring high-quality mappings that are critical to achieving high performance in constrained scenarios. This is especially relevant for low- to moderate-sized CGRA configurations, where the solution space, while large, remains within the tractable bounds for exact methods. A further benefit is the utilization of established solvers, such as Z3, which underscores the strategic integration of robust, industry-standard tools in our methodology.
@@ 4.1. Schedule creation
Table 2. Kernel Mobility Schedule.
@@ 4.2. SAT formulation
LABEL:eq:stm1 will be used on each node of the DFG and each $\beta$ generated will be added to the SAT solver. In our example with $n_{3}$ LABEL:eq:stm1 becomes:
@@ 4.3. Register Allocation
In our CFN formulation, to have a more straightforward set of constraints, we chose not to include information about which register should store the data produced by the node on every PE. We hence treat the RA task as a separate problem. Opting for a SAT formulation that yields register-oblivious solutions was a strategic decision aimed at minimizing compilation times. However, a consequence of this approach is the necessity to validate the solutions returned by the scheduling phase, in terms of register allocation needs. This issue is observable in only 3 of our experiments (backprop $3\times 3$, $4\times 4$, and $5\times 5$). Generally, after the solver returns a mapping, the RA phase needs to ensure that the CGRA has an adequate number of registers to execute the CIL correctly. SAT-MapIt, as a separate phase subsequent to SAT solving, exploits the SSA format of the input code and looks for an optimal solution(Hack and Goos, 2006) of the RA problem. For each PE in the CGRA, we extract all kind of dependencies referenced by that PE and we construct an interference graph. We then attempt to color using $n$ colors, where $n$ is the number of registers available on each PE.
If the coloring succeeds, we know that no further action is needed, and the toolchain provides the assembly code for the CGRA. However, if the coloring process fails, we increment the II and initiate a new search. While it’s possible to explore additional graph splitting techniques to make the graph colorable, we do not consider them herein, as such optimizations are orthogonal to the mapping problem we address in this work.
@@ 5. Experimental Results: SAT-MapIt solutions
In this section, we evaluate the effectiveness of SAT-MapIt on a set of CILs from MiBench and Rodinia benchmark suites, shown in Table 3. Table Table 3 provides more insights into the types of applications that are accelerated in our experiments. It also highlights how our methodology is applied to single loops, often focusing on the innermost ones, which were extracted utilizing the RAMP toolchain. The table includes information on the number of nodes, edges, and lines of code for each benchmark. It is important to note that these metrics alone do not fully capture the complexity involved in mapping a given loop, but they still enrich the understanding of the applications’ complexity.
We compare the obtained II, and the time to find it, with respect to two techniques of the SoA: RAMP(Dave et al., 2018) and PathSeeker (Balasubramanian and Shrivastava, 2022). We show results for only those benchmarks where RAMP and PathSeeker were able to provide a solution and where CFG was not present within the CIL body.
Table 3. List of benchmarks with graph information
Suite CIL #nodes #edges #lines of code
MiBench sha 30 33 1
sha2 26 28 1
gsm 20 24 3
patricia 42 46 3
bitcount 26 29 1
basicmath 19 20 6
stringsearch 16 16 2
Rodinia backpropagation 35 39 1
nw 16 16 1
srand 22 22 3
hotspot 67 76 8
@@ 5.2. SAT-MapIt achieves better II
It is worth noting that, in some benchmarks, expanding the array size fails to reduce the II (e.g. stringsearch). This arises from the presence of a long dependency chain within the DFG. Such chain acts as a bottleneck, limiting the potential for performance improvement through increased array size.
@@ 5.3. SAT-MapIt uses tight resources better
5.4. Runtime Analysis
Given that the II found are better than SoA, it may be interesting to also analyze the time needed to find such solutions; this is reported in Table 4. The following statements can be made when looking at these numbers. 1) Our tools running time is longer than SoA in 26 out of 44 experiments. 2) However, in these 26 cases the average time difference is only 15.28 seconds, with a standard deviation of $4.973$) On the other hand, in the 18 cases in which our tool is faster, the average time difference is 962.24 seconds, with a standard deviation of $1438.78$. This shows that, in these experiments, SAT-MapIt is significantly faster than state-of-the-art methods, particularly when the problem is challenging and hence computation times are high. The solver time will, of course, still increase worse-case exponentially with the size of the input, and it might not scale for some instances; however, where it does scale, it obtains a solution which is found on average tangibly faster than state of the art.
@@ 5.5. Limitations of SAT-MapIt
Currently, our tool does not apply any routing strategies. This limitation manifests in the sha CIL of a $5\times 5$ CGRA, where we achieve an II of 3, while (Dave et al., 2018) and (Balasubramanian and Shrivastava, 2022) can find an II of 2 by adding a routing node. This is the only case, out of the 44 experiments shown, where the effect of this limitation can be noticed. Implementing routing capabilities in SAT-MapIt will be the subject of future work. Another limitation can be seen in the optimality of SAT-MapIt, hence in its capability of returning only optimal solutions. Indeed, when an exact strategy does not scale (which is necessarily the case for some input instances, given the high complexity of the problem tackled) a non-exact algorithm could provide a non-guaranteed-optimal but good-enough feasible solution. An non-exact strategy for SAT-MapIt could involve restricting the solver’s exploration time for each II, hence potentially yielding sub-optimal yet feasible solutions. For instance, in our hotspot benchmark for a $2\times 2$ CGRA setup, no solutions emerged after a 4000 seconds (and even, in an additional experiment, after a 24-hour timeout). However, when we imposed an exploration limit of 5 minutes and 1 minute per II we detected a feasible solution at $II=21$ under both conditions. These adjustments led to total compilation times of 975.3 seconds with a 5-minute limit, and 257.2 seconds with a 1-minute limit. The exploration of non-exact strategies for SAT-MapIt is left to future work. A further limitation consists of the separation of the phases of scheduling and register allocation. The separation was a deliberate choice aimed at balancing the intricacy and computational efficiency of the problem-solving process. Our decision to decouple register allocation from the SAT formulation and address it in a subsequent phase is rooted in the desire to maintain clarity and manageability within the problem formulation. However, we acknowledge that this decoupling might lead to suboptimality, and this represents a potential direction for future work.
@@ 7.1. Experimental Setup
We consider square OpenEdgeCGRA instances with sizes $2\times 2$, $3\times 3$ and $4\times 4$, referred to as $D2$, $D3$ and $D4$, respectively, in the following. Benchmark CILs from the MiBench suite were executed on the architectures. To compare with the SoA in Section 5.2 we used the same DFGs generated by the RAMP toolchain; however those DFGs are not inherently compatible with the target CGRA’s ISA, preventing the direct execution of experiments using identical graphs. To address this discrepancy, we recalculated the DFGs for each CIL to align with the instruction set of OpenEdgeCGRA, for this part of the experiments.Table 6 shows the number of nodes and edges in the newly generated DFGs. Consequently, this resulted in different II values compared to Figure 7.
Table 6. List of benchmarks with graph information after the Instruction Selection phase targeting OpenEdgeCGRA
Benchmarks #nodes #edges
reversebits 9 10
bitcount 6 7
sqrt 8 12
stringsearch 16 18
gsm 14 20
sha 25 29
sha2 25 33
@@ 7.2. Experimental Results
[…]
Control intensive loops, or those with short kernels, pose a high strain on the CPU which needs to frequently perform counter increments, comparisons and jumps. On a CGRA, these tasks can be parallelized leading to speed-up of up to 3.8×3.8$\times$3.8 × among the benchmarked CILs. Other works have shown how CILs which make an intensive use of the memory can also greatly benefit from execution on the OpenEdgeCGRA (Carpentieri et al., 2024).
The sha kernel is the only case in which using the CGRA always translates into energy increase. Its access to discontiguous positions in memory requires the CGRA to recompute addresses on every iteration, demanding a large number of multiplications for which the ISA is not optimized.
One additional note-worthy observation from Table 8 is that the $D4$ CGRA is never the most energy-efficient, and that in some cases a trade-off between speed-up and energy exists (e.g. bitcount, sha). This underlines both the suitability of smaller CGRAs to accelerate edge kernels and the relevance of a translation between compiler-level and runtime metrics.
Table 8. Gains obtained from executing the benchmark CILs on the CGRA vs. executing them on the open-hardware X-HEEP platform from (Machetti et al., 2024).
Speed-up Energy gain
Kernel $D2$ $D3$ $D4$ $D2$ $D3$ $D4$
reversebits 3.8× 3.7× 3.6× 3.5× 2.5× 1.8×
bitcount 1.7× 1.6× 2.0× 1.8× 1.3× 1.3×
sqrt 1.7× 1.7× 1.6× 1.8× 1.4× 1.0×
stringsearch 2.3× 3.7× 3.7× 1.8× 2.3× 1.8×
gsm 1.4× 1.9× 1.7× 1.2× 1.2× 0.8×
sha 1.0× 1.4× 1.8× 0.7× 0.8× 0.7×
sha2 N/A 2.4× 2.4× N/A 1.3× 1.0×
Average 2.0× 2.4× 2.4× 1.8× 1.6× 1.2×